y= emx

5(dy/dx) = 2y


1
Expert's answer
2021-03-30T07:48:24-0400

Given that

"y = e^{mx}"

We have that

"\dfrac{dy}{dx} = me^{mx}"

Substituting into the equation, we have

"5me^{mx} = 2e^{mx}\\ 5me^{mx} -2e^{mx} = 0 \\ e^{mx}(5m-2) = 0"

Since "e^{mx} \neq 0" , we have that

"5m-2 = 0\\ 5m=2\\ m = \frac{2}{5} = 0.4"


Hence, "y = e^{0.4x}"